OpenStudy (anonymous):

3x^2-6=-153 solve the following equation using square root property

OpenStudy (angela210793):

3x^2=-153+6 3x^2=-147 x^2=-147/3 x^2=-49 x=sqr-49=7i

OpenStudy (anonymous):